OpenStudy (anonymous):

show work: Janine worked 5 1/2 hours on Saturday and 7 3/4 on Sunday. How many hours did she work on those two days.

OpenStudy (anonymous):

13 and 1/4 hours

OpenStudy (cwrw238):

1/2 + 3/4 = 2/4 + 3/4 = 5/4 or 1 1/4 5 + 7 + 1 1/4 = 13 1/4